A gas sample occupies a volume of 12.0 L at 50oC and 700 Torr. How many moles of gas are contained in the sample?


1
Expert's answer
2021-04-08T12:11:47-0400

V = 12 L

T = 50⁰C = 273 + 50 = 323 K

P = 700 torr

R = 62.363577 L.Torr / K.mol

PV = n RT

700×12 = n ×62.363577 × 323

n = 2.39 mol